Circular No. 9028

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 CL_{94} (LEMMON) An apparently asteroidal object discovered last year in the course of

the Mt. Lemmon survey (discovery observation tabulated below from MPS 237769) was re-discovered with the Mt. Lemmon 1.5-m reflector by S. M. Larson on 2009 Mar. 17.4 UT with an 8′′-diameter coma and a 15′′ diffuse tail in p.a. 310^{o} in four co-added images (and identified by T. Spahr). Four unfiltered 60-s co-added Mt. Lemmon CCD images taken by A. R. Gibbs on Mar. 18.3 in good seeing show a condensed coma of size 6′′ with a broad, 10′′ tail in p.a. 315^{o}. Following a request by B. G. Marsden, W. H. Ryan (Magdalena Ridge Observatory 2.4-m f/8.9 reflector) reports that his images from Mar. 18.2–18.4 also show a faint tail in p.a. ≈ 315^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
Feb. 8.39714  11^{h}27^{m}55.^{s}15  +11^{o}47′10′′.0  20.8

New astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2009-F28.

Epoch = 2006 July 4.0 TT

T = 2006 July 9.1877 TT  ω = 82.^{o}6364
e = 0.121042  Ω = 33.5426 }2000.0
q = 5.435878 AU  i = 8.3416

a = 6.184456 AU n^{o} = 0.0640843 P = 15.38 years

COMETS C/2008 W6–W10 (SOHO) Further to IAUC 9027, additional Kreutz sungrazers have been found

on SOHO website images, with their “discovery” observations tabulated below. C/2008 W6 was small and slightly diffuse (mag ∼ 7.5). C/2008 W7 was stellar in appearance (mag ∼ 7–7.5) in C3 images, and slightly diffuse with a faint, diffuse tail in C2 images. C/2008 W8 was tiny and stellar in appearance (mag ∼ 7.5). C/2008 W9 was diffuse (mag ∼ 7.5). C/2008 W10 was slightly diffuse (mag ∼ 7.5–8).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 W6 Nov. 24.504 16^{h}00.^{m}5 −22^{o}24′ C2  ZJ 2009-C56
C/2008 W7  24.779 15 58.1 −23 38 C3/2 JR 2009-C56
C/2008 W8  26.564 16 09.6 −22 47 C2  EB 2009-C57
C/2008 W9  27.629 16 13.7 −22 59 C2  RK 2009-C57
C/2008 W10  28.146 16 16.2 −23 03 C2  JR 2009-C57
2009 March 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green